OpenStudy (anonymous):

can some one give me this answer please x+4y=l6 _x+5y=11

OpenStudy (anonymous):

is that a -x instead of _X ?

OpenStudy (anonymous):

if it is then add the 2 equations like this to find y = ?

OpenStudy (anonymous):



OpenStudy (anonymous):

then substitute y=3 in the first equation to find x x + 12 = 16 x = 16-12 = 4

OpenStudy (anonymous):

do you understand the above? thanks

OpenStudy (blank ):

Already answered it. How could you have two open questions? http://openstudy.com/study#/updates/53895bb9e4b02fec9034b634

OpenStudy (anonymous):

damn!!! that's not fair ;) no wonder why no replies from the questioneer

OpenStudy (anonymous):

thanks Blank.